a 12kg uniform disc has a radius of 25cm and spins on an axis through it's centre and perpendicular to it's plane find the radius of it's gyration
1
Expert's answer
2020-10-20T07:11:03-0400

The moment of inertia of a disc of mass "M" and radius "R" about an axis through its center and perpendicular to its plane is given by

"I=\frac{1}{2}MR^2" ..............(1)

In terms of radius of gyration, "I=MK^2" ................(2)

Here, "I" is the moment of inertia and "K" is the radius of gyration.

From (1) and (2),

"MK^2=\frac{1}{2}MR^2\\
\Rightarrow K^2=R^2/2\\
\Rightarrow K=R/\sqrt 2"

The radius of the disc is "R=25\ cm"

Therefore, the radius of gyration is "K=\frac{25\ cm}{\sqrt 2} = 17.7\ cm"


Answer: The radius of gyration of the disc which spins on an axis through its centre and perpendicular to its plane is 17.7 cm.